OpenStudy (anonymous):

the um of two numbers re 145. the smaller number is decreased by 2 and the larger number is increased by 1. their ratio are 2:7. Find the two numbers ??? .. can you plss help me solving itt?? :((

OpenStudy (cwrw238):

let 2 numbers be x be smaller number and y larger then x + y = 145 also (x - 2) / (y + 1) = 2/7 solve for x and y

OpenStudy (cwrw238):

cross multiplying second equation: 7x - 14 = 2y + 2 7x - 2y = 16...............(1) multiply original first equation by 2: 2x + 2y = 290.............(2) now you can solve (1) and (2) by elimination method

OpenStudy (anonymous):

where 290 came from Friend???

OpenStudy (cwrw238):

145 * 2: multiplying x + y = 145 by 2 , so as to get -2y and adding will eliminate y

OpenStudy (cwrw238):

when multiply an equation you must multiply every term

OpenStudy (anonymous):

from the second equation friend, where do 2x came from?